OpenStudy (anonymous):

Write an equation of a line whose graph is parallel to the graph of y = 3x – 10.

OpenStudy (anonymous):

x=3 y= 5

jimthompson5910 (jim_thompson5910):

why did you write x = 3 and y = 5 ?

OpenStudy (anonymous):

y = 3x + b

OpenStudy (anonymous):

what i conclude is that the equation is parallel to 3

OpenStudy (anonymous):

and we let b the value of y-intercept

jimthompson5910 (jim_thompson5910):

are you supposed to find the equation of the line parallel to y = 3x - 10 and this parallel line goes through (3,5)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

is it the correct one

jimthompson5910 (jim_thompson5910):

y = mx + b y = 3x + b 5 = 3*3 + b 5 = 9 + b solve for b

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

b = -4

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the equation of the parallel line is y = 3x - 4

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

dont go i have a dilation multiple choices

jimthompson5910 (jim_thompson5910):

ok I'll help with 2 more

OpenStudy (anonymous):

are you going

jimthompson5910 (jim_thompson5910):

soon, but I can do 2 more

OpenStudy (anonymous):

mine are multiple choice can hou point the answe with the minimum explanation

jimthompson5910 (jim_thompson5910):

sure I'll see what I can do

OpenStudy (anonymous):

http://prntscr.com/zpm4e

OpenStudy (anonymous):

sorry A

jimthompson5910 (jim_thompson5910):

does it say if it's clockwise or counterclockwise?

OpenStudy (anonymous):

there is nothing written linke this

OpenStudy (anonymous):

but i suppose its counter clockwise

jimthompson5910 (jim_thompson5910):

if it's counter clockwise, then you add 240 degrees to 60 degrees to get 300 degrees

jimthompson5910 (jim_thompson5910):

which is where point K is at

OpenStudy (anonymous):

http://prntscr.com/zpm6n

OpenStudy (anonymous):

A is the correct option i think

jimthompson5910 (jim_thompson5910):

yes it's A

OpenStudy (anonymous):

http://prntscr.com/zpmi6

OpenStudy (anonymous):

i think this one is D

jimthompson5910 (jim_thompson5910):

no your first answer of C is correct

OpenStudy (anonymous):

but i got 0 in it

jimthompson5910 (jim_thompson5910):

oh, hmm ok let me recheck

jimthompson5910 (jim_thompson5910):

oh my bad, I'm thinking I and III

jimthompson5910 (jim_thompson5910):

so it would be B

OpenStudy (anonymous):

http://prntscr.com/zpmky

OpenStudy (anonymous):

D

OpenStudy (anonymous):

sory B

jimthompson5910 (jim_thompson5910):

C actually, it will help to plot the xy axis and the points of each step

OpenStudy (anonymous):

why not B

jimthompson5910 (jim_thompson5910):

if you draw an xy axis, plot the point (-2, -1) then reflect it over y = -5 you'll get (-2,-9)

jimthompson5910 (jim_thompson5910):

then reflect (-2, -9) over x = 1 to get (4, -9)

OpenStudy (anonymous):

ok thanks http://prntscr.com/zpmve

jimthompson5910 (jim_thompson5910):

for each point P, Q, and R, you swap x and y

OpenStudy (anonymous):

so B

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks a lot

jimthompson5910 (jim_thompson5910):

yw